KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)."

KyledaGreat:

\[\log _{_{3}}(x - 3) +\log _{3}(x + 6) = \log _{3}(5x - 10)\]

KyledaGreat:

If you wish to enter log or ln, you must use the keypad. If there is more than one solution, separate your answers with a comma. Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used. x = No Solution (∅)

KyledaGreat:

@vocaloid

KyledaGreat:

is the answer 4 , something else, or no solution ?

AZ:

You need to know some log properties: \( \log (a) + \log(b) = \log(a\cdot b)\) and \( \log (x) = \log(y)\) means that \( x = y\)

AZ:

\(\log _{_{3}}(x - 3) +\log _{3}(x + 6) = \log _{3}(5x - 10)\) First simplify the left side of the equation using the property I shared above. What do you get? Basically to combine the two logs being added, it's going to be the same as log of both terms being multiplied